A woman stands on a scale in a moving elevator. Her mass is 55.0 kg, and the combined mass of the elevator and scale is an additional 797 kg. Starting from rest, the elevator accelerates upward. During the acceleration, the hoisting cable applies a force of 9190 N. What does the scale read (in N) during the acceleration? Number Units
Answer #1

Balancing the forces,

(797 + 55) a = 9190 - (797 + 55) x 9.8

Acceleration, a = 0.986 m/s^2

Weight, W = 55 x (9.8 + 0.986) = 593.25 N

Comment in case any doubt please rate my answer....